When Boron 5 B 10 bombarded by neutron, alpha-particles are emitted.The resulting nucleus has the mass number? 11 7 6 15 DOP3 pg-15.48 Q43

When Boron 5B10 bombarded by neutron, alpha-particles are emitted.The resulting nucleus has the mass number?
 
  1. 11
  2. 7
  3. 6
  4. 15
DOP3 pg-15.48 Q43

Grade:12th pass

1 Answers

Harishwar IIT Roorkee
9 years ago
Hi Student,

The ans is 2

Boron5B10bombarded by 0n1neutron, 2He4 alpha-particles is emitted
So initial number of protons+ neutrons = 10 + 1=-11(neutron)
Final number of protons+ neutrons = 11-4 = 7 which is the mass number of resulting nucleus